1300. Watts enter a transformer with 900.0 Volts. If 1.500 A is needed how many volts will exit the transformer?



1
Expert's answer
2021-01-04T14:34:53-0500

"I_p\cdot U_p\approx I_s\cdot U_s\to P_p\approx1.5U_s\to U_s\approx P_p/1.5=1300/1.5=867(V)"